OpenStudy (anonymous):

A triangle and a parallelogram both have areas of 20 cm2 and bases of 5 cm. How do their heights compare?

OpenStudy (anonymous):

for the triangle A=(B*H)/2=>H=2A/B just plug in the values for the parm A=B*H=>H=A/B same as above then compare